OpenStudy (anonymous):

Ibuprofen is the active ingredient of Nurofen tablets. The half-life of ibuprofen is approximately 2 hours. A tablet contains 200 milligrams of ibuprofen and Joshua takes two tablets for a headache. 1. explain what is mean by the half-life of ibuprofen 2. what type of decay is represented by this context? 3.calculate the amount of ibuprofen in Joshua's system after 12 hours 4. draw a sketch graph to show how the amount of ibuprofen decrease over time 5. will this graph ever intersect the horizontal axis? explain your answer from a mathematical as well as a physiological perspective.

OpenStudy (anonymous):

1. The amount of time it takes for half of the quantity of a substance to disappear.

OpenStudy (anonymous):

2. exponential

OpenStudy (anonymous):

3. 12 hours = 6 half-lives 200/(2^6) milligrams left = 200/64 =3.125 milligrams

OpenStudy (anonymous):

4.